OpenStudy (anonymous):

Write an equation in point-slope form of the line through (2, 0) and (3, -2)

OpenStudy (anonymous):

m(x-x1)=y-y1 m=(-2-0)/(3-2)=-2 -2(x-2)=y y=-2x+4

OpenStudy (anonymous):

can you write that is slope-intercept form?

OpenStudy (anonymous):

y=mx+b m=-2,b=4

OpenStudy (anonymous):

Point slope formula: y-y1= m(x-x1) M is slope First find the slope using the slope formula: y2-y1/x2-x1 Do the math